The combustion of naphthalene (C_10 H_8), which releases 5150.1 kJ/mol, is often used to calibrate calorimeters. A 1.05 g sample of naphthalene is burned in a calorimeter, producing a temperature rise of 3.86 °C. Burning a 1.83 g sample of coal in the same calorimeter causes a temperature change of 4.90 °C. What is the energy density of the coal?
1
Expert's answer
2020-09-29T06:16:50-0400

Molar mass of naphthalene"=10\times 12+8=128\ g"

Moles in "1.05\ g" naphthalene"=\frac{1.05}{128}"

Let the total capacity of calorimeter per unit temperature change"=C"

Using energy balance,

"\frac{1.05}{128}\times 5150.5=C\times 3.86\implies C=10.946\ KJ/\degree C"

Now coal is burnt.

Let "S" be the energy density "(" Energy stored in unit mass ")" ,

Applying energy balance,

"1.83S=10.946\times 4.90\implies S=29.31\ KJ/g"